# Don’t Take the Premise for Granted: Evaluating the Premise Critique Ability of Large Language Models

### v1

@@ Abstract
Jinzhe Li1,3 Gengxu Li1  Yi Chang1,2,3111Corresponding authors  Yuan Wu1111Corresponding authors 1School of Artificial Intelligence, Jilin University 2Engineering Research Center of Knowledge-Driven Human-Machine Intelligence, MOE, China 3International Center of Future Science, Jilin University {lijz2121, lgx22}@mails.jlu.edu.cn, yichang@jlu.edu.cn, yuanwu@jlu.edu.cn
@@ 2.1 LLM Robustness Evaluation
Recent studies have established comprehensive benchmarks for evaluating LLM robustness against input perturbations, testing performance across vocabulary, syntactic, and semantic modifications Wang and Zhao (2024), character to semantic-level adversarial prompts Zhu et al. (2023), and punctuation noise in mathematical reasoning Abedin et al. (2025). Investigations into reasoning robustness have examined stability against adversarial noise and counterfactual statements Hoppe et al. (2025), the ability to identify logical fallacies in bilingual contexts Zhai et al. (2025), and chain-of-thought reasoning weaknesses through premature answers Xu et al. (2024), embedded trigger words Xiang et al. (2024), and modified reasoning steps Peng et al. (2024), with debate-style evaluations revealing limitations in maintaining correct reasoning against invalid arguments Wang et al. (2023a). However, these studies primarily focus on evaluating robustness against input perturbations and reasoning process attacks while overlooking models’ ability to actively identify and critically engage with erroneous premises in inputs.
@@ 2.2 False Premise Detection
Research addressing LLM responses to misinformation demonstrates persistent challenges in various contexts. Evaluations using real-world scenarios Guo et al. (2025b) and adversarial conditions Sakib et al. (2025) show models struggle with subtle misleading content despite varying resistance across architectures. Several approaches have been proposed to detect and mitigate false premises, such as retrieval-augmented logical reasoning frameworks Qin et al. (2025), attention mechanism constraints Yuan et al. (2024), presupposition verification techniques Kim et al. (2021); Yu et al. (2022), and specialized prompting methods Wang et al. (2023b). Recent work Fan et al. (2025) identifies a specific "MiP-Overthinking" phenomenon where reasoning-specialized LLMs produce verbose yet ineffective responses to ill-posed questions with missing premises. Despite these advances, most existing research centers on model behavior when faced with misinformation but overlooks logical inconsistencies or errors embedded in the prompts. In contrast, our work not only accounts for these issues but also conducts a comprehensive evaluation across multiple dimensions.
@@ 3.1 Definition of Premise Critique Ability
3.2 Overview of Data Construction
@@
  • Error Categories: We define four distinct categories of premise errors, each designed to evaluate a different aspect of a model’s ability to identify and reason about flawed inputs.
  • Difficulty: Problems span three levels of mathematical difficulty to evaluate model performance across varying levels of complexity, including Normal (sourced from the GSM8K Cobbe et al. (2021) dataset), Medium (Adapted from the Chinese College Entrance Examination portion within OlympiadBench He et al. (2024)), and Difficult (drawn from Omni-MATH Gao et al. (2024) problems with a difficulty rating above six).
  • Problem Variants: For each base flawed premise problem scenario constructed under a specific error category and difficulty, we generate three distinct problem variants for comparative evaluation. This includes the Original Problem which features correct premises and serves as a baseline, the Flawed Problem where an intentional error is introduced into the premises, and the Flawed Problem with Explicit Instruction, which is the flawed version augmented by a clear instruction that prompts the model to check for errors in the premises. The model’s successful critique of the Flawed Problem serves as a genuine reflection of its premise critique ability. In contrast, the introduction of the Flawed Problem with Explicit Instruction is intended to provide a comparative reference, illustrating how direct guidance influences error detection against the model’s inherent capabilities. This contrast reveals whether it relies on explicit prompts or demonstrates autonomous reasoning in premise evaluation, clarifying the basis of its analytical processes.
By constructing 100 base problems for each combination of error type and difficulty, we obtain 1200 unique base problems. Each is expanded into three distinct variants, yielding a total of 3600 problems. This structured and scalable design supports rigorous assessment of how premise critique ability is influenced by both error type and task complexity. Below, we elaborate on the four error categories that form the core of this evaluation framework, with further details on data construction procedures and representative examples provided in Appendix C.
@@ 3.3 Evaluation Metrics
PPCR APCR Normal Proactive Instructed PPCCR APCCR
@@
Table 1: Performance of both non-reasoning and reasoning LLMs on PCBench, Compareing the overall Premise Critique Rates (PPCR, APCR), along with the overall average Answer Lengths (where "Normal", "Proactive", and "Instructed" denote the model’s answers to the Original Problem, Flawed Problem, and Flawed Problem with Explicit Instruction respectively) and the corresponding Cost Ratios (PPCCR, APCCR) for the easy subset.Values that are underlined are considered the best-performing or most extreme among the models for each evaluation metric.
@@ 5 Further Discussion
As illustrated in Figure 4, the Deepseek-R1 model’s response to a contradictory problem reveals that the model did identify the flaw in the problem statement: the conflicting conditions where Maria used "1/4 of a skein" and "182 yards". The model’s internal thought process included repeated phrases such as "Wait, let me check again," reflecting a sustained effort to verify and reconcile these discrepant pieces of information, an effort marked by internal hesitation and cognitive conflict. Ultimately, the model chose to disregard the fractional information and proceeded with reasoning and calculations solely based on "182 yards", making a selective prioritization of problem-solving conditions on the user’s behalf.
Notwithstanding its recognition and engagement with the contradictory premises during the internal thought process, the model’s final answer failed to explicitly highlight this inconsistency. Instead of informing the user about the fundamental contradiction in the input or explaining the rationale behind choosing one condition over the other, the model directly provided a calculation result based on its selected condition. This behavior of "deciding on behalf of the user" while concealing the problem’s flaw, although enabling the production of a formal answer, severely undermines its reliability and utility as a "critical collaborator." Users remain unaware of the input’s problematic nature and cannot understand the model’s specific calculation path, which directly contradicts our definition of Premise Critique Ability, the capability that requires not only resistance to misdirection but also clear identification and explanation of input errors.
This gap between internal recognition and external articulation is not unique to a single instance but is observed across multiple reasoning models, such as o3-mini-medium and o4-mini-medium (cf. the illustrative case shown in Figure 5 in the appendix) which demonstrate abnormally low PPCR values despite achieving high APCR scores. This discrepancy might be attributed to the following phenomenon: while they successfully detect premise errors internally, their Premise Critique Ability is not outwardly demonstrated in their final responses.
This case with Deepseek-R1 simultaneously reflects the previous finding: flawed inputs deepen the reasoning model’s overthinking, as evidenced by its repeated internal deliberations over contradictory options. Despite expending computational resources on this internal processing, these models fail to incorporate their critical awareness into its final response.
@@ 6 Conclusion
While our study evaluates 15 representative LLMs, the generalizability of our findings, especially to newly emerged or untested models remains uncertain. Additionally, the restriction of our dataset to English and Chinese may overlook the influence of language itself, potentially limiting premise critique performance by not accounting for the diverse logical structures and cultural nuances across languages. Furthermore, PCBench’s focus on mathematical reasoning narrows its scope, as it does not consider the distinct reasoning patterns and premise types found in other natural sciences. Finally, our focus on textual data leaves multimodal contexts as a promising direction for future expansion. These limitations highlight opportunities to broaden premise critique research across models, languages, domains, and modalities.
@@ Appendix B More Cases
Figures 6 to 13 present eight examples that include both successful and failed cases across four types of errors, all generated by the Deepseek V3 model. These examples clearly illustrate the validity of our task design and its ability to distinguish different types of model behavior. They also provide evidence for the effectiveness of using LLMs as judges and demonstrate that our task can reliably assess a model’s Premise Critique Ability.
@@ C.1 Data Construction Pipeline
Corresponding detailed illustrations of construction pipeline are provided in Tables 3 to 7.
@@ C.2 Details of Dataset Statistics
[…]
@@ Appendix E Prompt Template
The prompt template presents in Table 8 is designed for premise extraction. Prompts across Tables 9 to 13 are utilized to synthesize four types of errors, while the prompt in Table 14 is employed to evaluate whether models exhibit premise critique ability.
Figure 5: An illustrative case of o4-mini’s response to a Contradictory Inference Insertion Question. The question presented in the image contains a logical contradiction (highlighted in red). According to the statement that "the number of students doubles at the end of each month," the student population should undergo five doublings by the end of May. However, the question also claims that the final number of students is obtained by multiplying the initial amount by $2^{3}$, which implies only three doublings. This discrepancy between the implied five doublings and the stated three is the source of the contradiction. During the reasoning process, the model (highlighted in orange) notes that "the key part mentions multiplying by 2 raised to the power of 3, so we’ll focus on the population growth up to that 3rd power, and there may be a mix-up between doubling across 5 months and what the question asks," demonstrating that the model has identified the inconsistency regarding the number of doublings. However, in its final answer, the model does not further explore this insight. Instead, it proceeds with the calculation using $2^{3}$ and concludes that the number of additional students is 70.
@@
Part Of the Solution: First, find out the number of apples left after giving some to her brother. The number of remaining apples is $15-3=12$. Then, to find out how many apples each friend gets, we divide the number of friends by the remaining apples.
@@
"modified_part_solution":"First, find out the number of apples left after giving some to her brother. The number of remaining apples is $13-3=10$. Then, to find out how many apples each friend gets, we divide the number of friends by the remaining apples."
"correct_step":"The number of remaining apples is $15-3=12$."
"error_step":"The number of remaining apples is $13-3=10$."
@@
Correct Solution: First, find out the number of apples left after giving some to her brother. The number of remaining apples is $15-3=12$. Then, to find out how many apples each friend gets, we divide the number of friends by the remaining apples…
@@
"modified_wrong_solution":"First, find out the number of apples left after giving some to her brother. The number of remaining apples is $13-3=10$. Then, to find out how many apples each friend gets, we divide the number of friends by the remaining apples..."
@@
"correct_step":"The number of remaining apples is $15-3=12$."
"error_step":"The number of remaining apples is $13-3=10$."

### v2

@@ Abstract
Jinzhe Li1,3222These authors contributed equally.   Gengxu Li1222These authors contributed equally.   Yi Chang1,2,3111Corresponding authors   Yuan Wu1111Corresponding authors 1School of Artificial Intelligence, Jilin University 2Engineering Research Center of Knowledge-Driven Human-Machine Intelligence, MOE, China 3International Center of Future Science, Jilin University {lijz2121, lgx22}@mails.jlu.edu.cn, yichang@jlu.edu.cn, yuanwu@jlu.edu.cn
@@ 1 Introduction
Benchmark Ability Error Category Count Proactive Overthink Difficulty Comparison
ECHOMIST Guo et al. (2025b) Implicit misinfo handling 5 categories × ×
MathClean Liang et al. (2025) Math data validation 7 categories × × ×
MathTrap Zhao et al. (2024) Compositional generalization 5 categories × ×
RuozhiBench Zhai et al. (2025) Logical fallacy identification 6 categories × × × ×
MiP dataset Fan et al. (2025) Missing premise detection 4 subcategories (under 1 major category) ×
PCBench (Ours) Premise critique 4 categories
Table 1: Comparison between PCBench and other benchmarks. "Ability" describes the core capability under evaluation; "Category Count" indicates the number of error categories/subcategories of the benchmark; "Proactive" indicates prompting the model without any hints; "Overthink" denotes the consideration of overthinking phenomena; "Difficulty" represents difficulty grading; "Comparison" means including the comparison of the model’s performance under proactive (without explicit instruction) and passive (with explicit instruction) modes.
@@ 2.1 LLM Robustness Evaluation
Recent studies have established comprehensive benchmarks for evaluating LLM robustness against input perturbations, testing performance across vocabulary, syntactic, and semantic modifications Wang and Zhao (2024), character to semantic-level adversarial prompts Zhu et al. (2023), and punctuation noise in mathematical reasoning Abedin et al. (2025). Investigations into reasoning robustness have examined stability against adversarial noise and counterfactual statements Hoppe et al. (2025), the ability to identify logical fallacies in bilingual contexts Zhai et al. (2025), and chain-of-thought reasoning weaknesses through premature answers Xu et al. (2024), embedded trigger words Xiang et al. (2024), and modified reasoning steps Peng et al. (2024), with debate-style evaluations revealing limitations in maintaining correct reasoning against invalid arguments Wang et al. (2023). However, these studies primarily focus on evaluating robustness against input perturbations and reasoning process attacks while overlooking models’ ability to actively identify and critically engage with erroneous premises in inputs.
@@ 2.2 False Premise Detection
As shown in Table 1, multiple benchmarks evaluate LLMs’ handling of flawed premises. RuozhiBench Zhai et al. (2025) tests their capacity to detect logical fallacies and misleading premises via content from a Chinese forum with brain teasers. MathClean Liang et al. (2025) evaluates their capability to judge math problem/answer correctness and pinpoint errors. MathTrap Zhao et al. (2024) examines their compositional generalization in math reasoning through "trap problems". Among recent efforts, ECHOMIST Guo et al. (2025b) assesses LLMs’ ability to address implicit misinformation using real dialogues, social media data and rewritten queries. The MiP dataset Fan et al. (2025) identifies the "MiP-Overthinking" phenomenon: LLMs generate verbose yet ineffective responses to ill-posed questions with missing premises, preliminarily revealing model behavior under this unprompted flawed input. Building on these, our work extends to more flawed input types, conducts multi-perspective evaluation, and compares model critique performance between active and passive modes, better revealing inadequacies in LLMs’ proactive premise critique ability.
@@ 3.1 Definition of Premise Critique Ability
Figure 2: An overview of the dataset construction and the evaluation pipeline.
3.2 Data Construction
@@
3.2.1 Difficulty
Problems span three levels of mathematical difficulty to evaluate model performance across varying levels of complexity, including Normal (sourced from the GSM8K Cobbe et al. (2021) dataset), Medium (Adapted from the Chinese College Entrance Examination portion within OlympiadBench He et al. (2024)), and Difficult (drawn from Omni-MATH Gao et al. (2024) problems with a difficulty rating above six).
3.2.2 Error Categories
We define four distinct categories of premise errors, each designed to evaluate a different aspect of a model’s ability to identify and reason about flawed inputs:
@@ 3.2.2 Error Categories
3.2.3 Problem Variants
For each base flawed premise problem scenario constructed under a specific error category and difficulty, we generate three distinct problem variants for comparative evaluation. This includes the Original Problem which features correct premises and serves as a baseline, the Flawed Problem where an intentional error is introduced into the premises, and the Flawed Problem with Explicit Instruction, which is the flawed version augmented by a clear instruction that prompts the model to check for errors in the premises. The model’s successful critique of the Flawed Problem serves as a genuine reflection of its premise critique ability. In contrast, the introduction of the Flawed Problem with Explicit Instruction is intended to provide a comparative reference, illustrating how direct guidance influences error detection against the model’s inherent capabilities. This contrast reveals whether it relies on explicit prompts or demonstrates autonomous reasoning in premise evaluation, clarifying the basis of its analytical processes.
3.2.4 Construction Pipeline
As illustrated in Figure 2, the PCBench construction pipeline consists of four main steps. First, problems of Normal, Medium, and Hard difficulty levels are randomly sampled. Then, necessary premises are extracted from the sampled problems and then rewritten into fake premises to form synthesized false premises. Third, flawed problems with four types of erroneous premises are synthesized, covering Contradictory Premise Insertion, Contradictory Inference Insertion, Flawed Solution Completion, and Irrelevant Query Distraction. Fourth, human verification is conducted to screen out eligible flawed problems whose error types conform to the defined criteria. Finally, each verified problem is expanded into three variants: the Original Problem (serving as a baseline), the Flawed Problem with explicit instruction, and the Flawed Problem without explicit instruction. In total, we construct 100 base problems for each combination of the 4 error types and 3 difficulty levels, resulting in 1200 unique base problems with each base problem expanded into three distinct variants. This structured and scalable design enables rigorous assessment of how premise critique ability is influenced by both error type and task complexity.
@@ 3.3 Evaluation Metrics
PPCR APCR Original Proactive Instructed PPCCR APCCR
@@
Table 2: Performance of both non-reasoning and reasoning LLMs on PCBench, Compareing the overall Premise Critique Rates (PPCR, APCR), along with the overall average Answer Lengths (where "Original", "Proactive", and "Instructed" denote the model’s answers to the Original Problem, Flawed Problem, and Flawed Problem with Explicit Instruction respectively) and the corresponding Cost Ratios (PPCCR, APCCR) for the easy subset.Values that are underlined are considered the best-performing or most extreme among the models for each evaluation metric.
@@ 4.2 Main Results
Human Verification
To verify alignment between the LLM evaluator (o3-mini-high) and human evaluations, we conducted a validation study: 240 instances (60 per error type, 20% of each group) were randomly selected from premise critique results of GPT-4.1 and Qwen3-235B-A22B-thinking. Graduate students applied the same criteria as o3-mini-high (labeling responses as "True" if they clearly questioned premises, identified contradictions, and expounded on errors; otherwise "False").
Confusion matrices for the two models are shown in Table 16 (Qwen3-235B-A22B-thinking) and Table 17 (GPT-4.1). Kappa coefficients, 0.8921 for Qwen3-235B-A22B-thinking and 0.7763 for GPT-4.1, indicate strong consistency between automatic and human evaluations, confirming the model’s reliability as an automatic evaluator and the validity of our results.
@@ 5 Further Discussion
As shown in Figure 5, Deepseek-R1 identifies the contradiction in the problem: Maria’s use of "1/4 of a skein" and "182 yards" as conflicting conditions. Its internal reasoning (e.g., repeated "Wait, let me check again") reflects hesitation and efforts to reconcile the discrepancy, but it ultimately ignores the fractional information and calculates solely based on "182 yards," prioritizing conditions without user input.
Despite recognizing the contradiction internally, the model’s final answer does not explicitly highlight this inconsistency. It provides a result based on its chosen condition instead of informing users of the input flaw or explaining its reasoning. This "decision-making on the user’s behalf" undermines its reliability as a "critical collaborator," as users remain unaware of the problem’s issues and the model’s calculation path, contradicting our definition of Premise Critique Ability—the capability that requires not only resistance to misdirection but also clear identification and explanation of input errors.
This gap between internal recognition and external expression is not unique to a single instance but is observed across multiple reasoning models: models like o3-mini-medium and o4-mini-medium (see Figure 6 in the appendix) also show low PPCR despite high APCR, likely because they detect premise errors internally but fail to demonstrate Premise Critique Ability in final responses.
The case also supports the earlier finding that flawed inputs deepen models’ overthinking (evidenced by Deepseek-R1’s repeated deliberation). Despite expending computational resources on this internal processing, these models fail to incorporate their critical awareness into their final response.
@@ 6 Conclusion
Acknowledgments
The authors would like to thank the anonymous reviewers for their valuable comments. This work is supported by the National Key Research and Development Program of China (No.2023YFF0905400), the National Natural Science Foundation of China (No.U2341229) and the Reform Commission Foundation of Jilin Province (No.2024C003).
@@ Acknowledgments
As a preliminary field study, though we evaluated 15 representative LLMs, the generalizability of our findings, especially to new or untested models, remains uncertain. While our constructed flawed data partially aligns with real-world scenarios, a comprehensive, systematic statistical analysis of flaw probability and scenario classification in real-world user inputs is lacking. Meanwhile, using existing benchmarks may raise data contamination concerns, requiring deeper analysis. Additionally, limiting our dataset to English and Chinese may overlook language’s influence, potentially restricting premise critique performance by ignoring diverse logical structures and cultural nuances across languages. Furthermore, PCBench’s focus on mathematical reasoning narrows its scope, failing to consider distinct reasoning patterns and premise types in other natural sciences. Finally, our focus on textual data leaves multimodal contexts as a promising future direction. These limitations point to opportunities to expand premise critique research across models, languages, domains, and modalities.
@@ Appendix B More Cases
Figures 7 to 14 present eight examples that include both successful and failed cases across four types of errors, all generated by the Deepseek V3 model. These examples clearly illustrate the validity of our task design and its ability to distinguish different types of model behavior. They also provide evidence for the effectiveness of using LLMs as judges and demonstrate that our task can reliably assess a model’s Premise Critique Ability.
@@ C.1 Data Construction Pipeline
Corresponding detailed illustrations of construction pipeline are provided in Tables 4 to 8.
@@ C.2 Details of Dataset Statistics
The dataset is constructed by combining four distinct error categories with three levels of mathematical difficulty. For each unique combination of error category and difficulty level, 100 base problems are created. This results in a total of $4\text{ (Error Categories)}\times 3\text{ (Difficulties)}\times 100\text{ (Base Problems per Combination)}=1200$ unique base problems, as shown in Table 3.
Furthermore, each of these 1200 base problems is expanded into three distinct variants: the Original Problem (with correct premises), the Flawed Problem (with an intentional error in the premises), and the Flawed Problem with Explicit Instruction (the flawed version appended with the explicit prompt: "Check if there are any errors in the question’s premises before answering. If there are, please report them promptly."). This expansion yields a total dataset size of $1200\text{ (Base Problems)}\times 3\text{ (Problem Variants)}=3600$ problems. This structured design enables a thorough assessment of LLMs’ premise critique capabilities across diverse error types, varying difficulty tiers, and different prompting scenarios.
@@ Appendix E Prompt Template
The prompt template presents in Table 9 is designed for premise extraction. Prompts across Tables 10 to 14 are utilized to synthesize four types of errors, while the prompt in Table 15 is employed to evaluate whether models exhibit premise critique ability.
Appendix F More Results
Figures 15 to 17 illustrates more results of the evaluated LLMs.
Figure 6: An illustrative case of o4-mini’s response to a Contradictory Inference Insertion Question. The question presented in the image contains a logical contradiction (highlighted in red). According to the statement that "the number of students doubles at the end of each month," the student population should undergo five doublings by the end of May. However, the question also claims that the final number of students is obtained by multiplying the initial amount by $2^{3}$, which implies only three doublings. This discrepancy between the implied five doublings and the stated three is the source of the contradiction. During the reasoning process, the model (highlighted in orange) notes that "the key part mentions multiplying by 2 raised to the power of 3, so we’$l$ focus on the population growth up to that 3rd power, and there may be a mix-up between doubling across 5 months and what the question asks," demonstrating that the model has identified the inconsistency regarding the number of doublings. However, in its final answer, the model does not further explore this insight. Instead, it proceeds with the calculation using $2^{3}$ and concludes that the number of additional students is 70.
@@ Appendix F More Results
Figure 15: This heatmap illustrates the Proactive Premise Critique Rate (PPCR) of all the evaluated LLMs across different difficulty levels (normal, medium, hard) in PCBench.
Figure 16: This heatmap depicting the Proactive Premise Critique Rate (PPCR) of all the evaluated LLMs across four error categories (Contradictory Inference Insertion, Contradictory Premise Insertion, Flawed Solution Completion, Irrelevant Query Distraction) in PCBench.
Figure 17: This dot plot compares the Proactive Premise Critique Rate (PPCR, blue dots) and Assisted Premise Critique Rate (APCR, red dots) of all the evaluated LLMs. It reveals a significant gap between the models’ proactive premise critique ability (autonomously detecting flawed premises without explicit prompting) and their passive critique ability under prompting (detecting flawed premises with explicit instructions to check for errors), indicating most models heavily rely on explicit prompts for effective flawed premise identification.
@@
Part Of the Solution: First, find out the number of apples left after giving some to her brother. The number of remaining apples is 153=1215-3=12. Then, to find out how many apples each friend gets, we divide the number of friends by the remaining apples.
@@
"modified_part_solution":"First, find out the number of apples left after giving some to her brother. The number of remaining apples is 133=1013-3=10. Then, to find out how many apples each friend gets, we divide the number of friends by the remaining apples."
"correct_step":"The number of remaining apples is 153=1215-3=12."
"error_step":"The number of remaining apples is 133=1013-3=10."
@@
Correct Solution: First, find out the number of apples left after giving some to her brother. The number of remaining apples is 153=1215-3=12. Then, to find out how many apples each friend gets, we divide the number of friends by the remaining apples…
@@
"modified_wrong_solution":"First, find out the number of apples left after giving some to her brother. The number of remaining apples is 133=1013-3=10. Then, to find out how many apples each friend gets, we divide the number of friends by the remaining apples..."
@@
"correct_step":"The number of remaining apples is 153=1215-3=12."
"error_step":"The number of remaining apples is 133=1013-3=10."
@@
Confusion Matrix Human Prediction
Positive Negative
Model Prediction Positive 147 0
Negative 12 81
Table 16: Confusion Matrix for Qwen3-235B (Model vs. Human), with a Corresponding Kappa Coefficient of 0.8921
Confusion Matrix Human Prediction
Positive Negative
Model Prediction Positive 166 3
Negative 18 53
Table 17: Confusion Matrix for GPT-4.1 (Model vs. Human), with a Corresponding Kappa Coefficient of 0.7763